### v1

# Stable-Hair: Real-World Hair Transfer via Diffusion Model

Abstract
Current hair transfer methods struggle to handle diverse and intricate hairstyles, thus limiting their applicability in real-world scenarios. In this paper, we propose a novel diffusion-based hair transfer framework, named Stable-Hair, which robustly transfers a wide range of real-world hairstyles onto user-provided faces for virtual hair try-on. To achieve this goal, our Stable-Hair framework is designed as a two-stage pipeline. In the first stage, we train a Bald Converter alongside stable diffusion to remove hair from the user-provided face images, resulting in bald images. In the second stage, we specifically designed three modules: a Hair Extractor, a Latent IdentityNet, and Hair Cross-Attention Layers to transfer the target hairstyle with highly detailed and high-fidelity to the bald image. Specifically, the Hair Extractor is trained to encode reference images with the desired hairstyles. To preserve the consistency of identity content and background between the source images and the transfer results, we employ a Latent IdentityNet to encode the source images. With the assistance of our Hair Cross-Attention Layers in the U-Net, we can accurately and precisely transfer the highly detailed and high-fidelity hairstyle to the bald image. Extensive experiments have demonstrated that our approach delivers state-of-the-art (SOTA) results among existing hair transfer methods. Project page: https://xiaojiu-z.github.io/Stable-Hair.github.io/
Figure 1: Stable-Hair is a novel diffusion-based hairstyle transfer method that can robustly transfer a diverse range of real-world hairstyles. We demonstrate its performance on a variety of challenging hairstyles, achieving highly detailed and high-fidelity transfers with impressive results while preserving the original identity content and structure.
Introduction
With the increasing prevalence of digital media and virtual reality applications, personalized virtual avatars and virtual try-on systems have emerged as a significant research area. Hair transfer is one of the most challenging tasks within this domain. The objective of this task is to transfer the hair color, shape, and structure attributes from a reference image to a user-provided source image while preserving the identity and background of the source image.
In recent years, the advancements in Generative Adversarial Networks (GANs) (Tan et al. 2020; Zhang and Zheng 2018; Guo et al. 2022; Zhu et al. 2022; Chang, Kim, and Kim ; Chung et al. 2022; Zhu et al. 2021; Saha et al. 2021; Wei et al. 2022, 2023; Nikolaev et al. 2024; Khwanmuang et al. 2023; Kim et al. 2022) have driven significant progress in this field. However, these GAN-based methods often struggle to handle the diverse and complex hairstyles encountered in real-world scenarios, which severely limits their effectiveness in practical applications. Recently, diffusion models have emerged as state-of-the-art methods in the field of image generation. These models not only enable more stable training but also demonstrate impressive results in terms of diversity and fidelity. Consequently, we are intrigued by the question: “Can we leverage the powerful capabilities of diffusion models to achieve more stable and high-precision hair transfer? ”
Building on the aforementioned considerations, in this paper, we propose a novel hair transfer framework based on diffusion models, named Stable-Hair. Inspired by method (Wei et al. 2023) which converts different editing conditions (e.g., text, sketch) into different proxies in the StyleGan W+ space, we designed a two-stage paradigm to ensure the precision and naturalness of our hair transfer process. In the first stage, we train a bald converter to transform the user-provided source image into a bald proxy image. In the second stage, we utilize our specially designed model to transfer the reference hairstyle onto the bald proxy image.
Our model is designed with three core components, each addressing specific aspects of the hair transfer challenge: the Hair Extractor, Latent IdentityNet, and Hair Cross-Attention Layers. The Hair Extractor is specially trained to capture and encode hairstyles from reference images with unprecedented levels of detail and texture. This capability enables Stable-Hair to accurately represent complex hairstyles. The Latent IdentityNet is used to encode the source image, ensuring that the identity and background of the source image remain consistent throughout the transfer process. By integrating Hair Cross-Attention Layers within diffusion U-Net, we can effectively inject the features of the reference hair into the source image. This approach not only allows the hair to adapt accurately to the new environment of the source image but also ensures that it is styled appropriately, appearing natural and matching the original appearance of the subject.
Notably, to preserve non-hair regions and the overall facial identity of the subject, we have designed a novel structure called Latent Controlnet. This structure integrates our Bald Converter and Latent IdentityNet. By transforming the pixel-space transfer process to operate within the latent space of the diffusion model, Latent Controlnet ensures color consistency in the non-hair regions of the source image during the transfer process.
Through extensive experiments, Stable-Hair has demonstrated its superior performance, significantly surpassing existing state-of-the-art hair transfer methods in terms of fidelity and fine-grained detail. Our approach sets a new standard for hair transfer technology, promising to revolutionize the virtual hair try-on experience with its advanced capabilities and innovative design.
In summary, our contributions are:
  • In this paper, we propose the first diffusion-based hair style transfer framework, named Stable-Hair. Compared with the GAN-based hair style transfer works, Stable-Hair achieves better fidelity and robustness. Experiments demonstrate that our method surpasses other state-of-the-art methods.
  • We utilize a Hair Extractor in conjunction with Hair Cross-attention Layers to inject reference hair features. To ensure better source content consistency during the transfer process, we introduce a novel Latent ControlNet architecture. This architecture is utilized as a Bald Converter and a Latent IdentityNet, which maps the hairstyle transfer process from the pixel space to the latent space.
  • We devised an automated data production pipeline, which is a pivotal component contributing to the successful training of our framework.
Related Works
Hair style transfer.
The development of GAN-based methods (Tan et al. 2020; Zhang and Zheng 2018; Guo et al. 2022; Zhu et al. 2022; Chang, Kim, and Kim ; Chung et al. 2022; Zhu et al. 2021; Saha et al. 2021; Wei et al. 2022, 2023; Nikolaev et al. 2024; Khwanmuang et al. 2023; Kim et al. 2022; Shu et al. 2022) has significantly advanced the field of hairstyle transfer, with most current methods relying on GAN-based approaches.
In the realm of hair transfer, MichiGAN(Tan et al. 2020) decomposes hair into four orthogonal attributes and designs corresponding modules to represent, process, and convert user inputs, which are then integrated to realize an end-to-end network. Similarly, Barbershop(Zhu et al. 2021) proposes a novel latent space for image blending that excels in preserving detail and encoding spatial information, extracting hairstyle information from multiple reference images. Furthermore, LOHO(Saha et al. 2021) employs an optimization-based approach using GAN inversion to infill missing hair structure details in the latent space during hairstyle transfer, introducing two-stage optimization and gradient orthogonalization to enable disentangled latent space optimization of hair attributes. Moreover, Hairmapper(Wu, Yang, and Jin 2022) focuses on training a fully connected network to determine the direction of hair removal in the latent space of StyleGAN during the training stage. SYH(Kim et al. 2022) explores a pose-aligned latent code in StyleGAN2’s latent space and proposes a high-performing pose-invariant hairstyle transfer model, which excels in transferring hairstyles under significant pose differences while preserving local hairstyle textures. In addition, Hairclip(Wei et al. 2022), leveraging the powerful image-text representation capability of the Contrastive Language-Image Pre-Training (CLIP) model, proposes a unified hair editing framework based on StyleGAN. Building on this, HairCLIPv2(Wei et al. 2023) converts all hair editing tasks into hair transfer tasks, with editing conditions converted into different proxies accordingly. It also supports a wide range of interaction modes and functions. StyleGAN-Salon(Khwanmuang et al. 2023) introduces a multi-view optimization framework and the use of two guide images, representing a rough composite of the final output based on a simple cut-and-paste from the input image’s viewpoint, to optimize StyleGAN2’s learned latent space on W and W+, thereby controlling hallucinations to accurately transfer hairstyles. HairFastGAN(Nikolaev et al. 2024) addresses the challenge of hairstyle transfers in different poses and proposes a new encoder capable of generating images promptly.
Despite these advancements, GAN-based models have limitations in addressing complex hairstyle transfers in real-world scenarios. To overcome these deficiencies, we propose the first diffusion-based method, which generates high-quality and robust images, thereby setting a new benchmark in the field.
Diffusion models.
Currently, diffusion models have garnered widespread attention and have seen significant advancements. As the most prominent generative models today, diffusion models have achieved state-of-the-art results across various image generation tasks, including text-to-image generation (Ramesh et al. 2022; Podell et al. 2023; Saharia et al. 2022a; Rombach et al. 2022; Saharia et al. 2022b), image editing (Chefer et al. 2023; Li et al. 2023; Epstein et al. 2023; Cao et al. 2023; Xie et al. 2023; Mou et al. 2023a; Bodur et al. 2023; Zhang et al. 2023; Tsaban and Passos 2023; Yang et al. 2024; Li et al. 2024), controllable generation (Zhang and Agrawala 2023; Mou et al. 2023b; Zhao et al. 2023; Ma et al. 2023), personalized image generation (Zhang et al. 2024b; Hu et al. 2022; Arar et al. 2023; Jia et al. 2023; Gal et al. 2022; Ruiz et al. 2023, 2022; Zhang et al. 2024c) and so on. Notably, the recent emergence of diffusion-based virtual try-on applications further demonstrates the formidable generative capabilities of diffusion models, enabling commercial-grade virtual try-ons (Xu et al. 2024; Wang et al. 2024; Kim et al. 2023; Zeng et al. 2023) and virtual makeovers (Zhang et al. 2024d), which were previously unattainable with traditional GAN methods.
Diffusion models exhibit a pronounced advantage in generating high-quality images, especially when trained on extensive and diverse datasets. In this paper, we harness the strengths of diffusion models to produce high-fidelity and robust output images. This approach is particularly effective in overcoming the limitations inherent in GAN-based models, especially in the context of complex hairstyle transfer tasks. By leveraging the superior generative capabilities of diffusion models, our method significantly enhances the quality and realism of the transferred hairstyles, thereby addressing the challenges posed by intricate and detailed hair structures.
Methodology
Stable-Hair
Overview.
Figure 2: Comparative structural diagram of ControlNet and Latent ControlNet (ours).
Our design divides the hairstyle transfer process into two stages. Firstly, the user’s input source image is transformed into a bald proxy image by using the Bald Converter. Secondly, our model is employed to transfer the reference hair onto the bald proxy image. This ensures optimal stability in hairstyle transfer and maintains the source image content.
Figure 3: Overall schematics of our method. Our pipeline consists of two stages to achieve high-quality hair transfer. First, the user’s input source image is transformed into a bald proxy image. This transformation is accomplished using a pre-trained Stable Diffusion (SD) model in conjunction with a specialized Bald Converter. In the second stage, we employ the pre-trained SD model along with a Hair Extractor to transfer the reference hair onto the bald proxy image. The Hair Extractor is responsible for capturing the intricate details and features of the reference hair. These features are then injected into the SD model through newly added hair cross-attention layers. By leveraging these two stages, our method achieves highly detailed and high-fidelity hair transfers, producing natural and visually appealing results.
As shown in Fig 3, our Stable-Hair framework comprises three primary components: Hair Extractor, Latent IdentityNet, and the newly integrated Hair Cross-Attention Layers. Given a reference image $\mathit{I_{r}}$ containing the desired hairstyle and a source image $\mathit{I_{s}}$ provided by users, Hair Extractor can stably and precisely encode a variety of reference hairstyles in real-world scenes and extract detailed hair features $\mathit{F_{h}}$. These detailed hair features $\mathit{F_{h}}$ are fed to the diffusion U-Net structure through Hair Cross-Attention Layers. The Latent IdentityNet modules are employed for processing bald proxy image $\mathit{I_{b}}$ which is converted from the source image $\mathit{I_{s}}$ by Bald Converter.
Figure 4: Accumulated color deviations in different designing options.
Latent ControlNet.
In our method, the source image is processed in two steps, The first step is to convert the source image into a bald proxy image by a Bald Converter. The second step is to maintain the content of the non-hair region of the bald proxy image using Latent IdentityNet. The maintenance of content consistency in the source image is crucial through these two steps. Any deviation from the intended goal, such as alterations in color or structure, would result in a content-inconsistent final image. Therefore, designing these two content maintenance modules becomes a pivotal aspect of our hair transfer framework.
A naive baseline approach involves utilizing the ControlNet structure as a Bald Converter and Latent IdentityNet to ensure content consistency. However, our experimental findings indicate that while ControlNet effectively maintains the structural consistency of the source image, it struggles to preserve color consistency. As shown in Fig. 4, due to accumulated color deviations over these two steps, there are noticeable changes in final colors. Why does the controlnet have a color difference? We supposed that the reason is that pixel space and latent space represent image information in fundamentally different ways. Pixel space deals with the raw pixel values of an image, while latent space involves a more abstract, high-dimensional representation created by VAE encoder. For diffusion models, the characteristics and distribution of information in these two spaces can vary significantly, making alignment challenging during training.
Therefore, as depicted in Fig. 2, we improve the ControlNet structure and propose a new variant named Latent ControlNet. Before the image is input into the ControlNet, the image is encoded into the latent space by the VAE encoder and then sent to the trainable copy of U-Net after a new trainable convolutional layer. Finally, we train our Bald Converter and Latent IdentityNet based on our proposed Latent ControlNet structure and get the best consistency practice.
Hair Extractor.
Hair transfer needs to transfer the hair on the reference image $\mathit{I_{r}}$ in a detailed, complete, and accurate way, and our Hair Extraxtor is designed to achieve such transfer. Inspired by recent works (Wang et al. 2024; Zhang et al. 2024a) on reference image-guided generation, we utilize a trainable copy of U-Net architecture from pre-trained diffusion models as our Hair Extractor. Specifically, we encode our reference image $\mathit{I_{r}}$ through Hair Extractor and collect the features of the self-attention layer in each transformer block as detailed hair features $\mathit{F_{h}}$. Then the features are subsequently injected into diffusion U-Net through Hair Cross-Attention Layers.
Hair Cross-Attention Layers.
The goal of the Hair Cross-Attention Layers is to integrate detailed hair features into the U-Net and align them appropriately with the intermediate feature maps of U-Net, enabling the model to accurately transfer the hair onto the appropriate region in the source image. Specifically, as shown in the figure, in each transformer block of the U-Net, we retain the original self-attention layers and add Hair Cross-Attention Layers. The detailed Hair Features $\mathit{F_{h}}$ are input to the hair cross-attention layers and serve as the K(key) and V(value) features. The hair cross-attention layers and the original self-attention layers in the U-Net share the Q feature. With the help of the cross-attention mechanism, the hair and source face are well-aligned. Finally, we simply add the output of hair cross-attention to the output of original self-attention.
Figure 5: Synthetic Training Data: We propose an automated data generation pipeline to generate (Original image, Reference image, Bald proxy image) triplets for training. The pipeline uses ChatGPT to generate text prompts, the Stable Diffusion Inpainting model to generate reference images, and our pre-trained Bald converter to convert the original image into the bald proxy image.
Model Training and Inference
Training Data Collection.
As shown in Fig. 5 Our pipeline for creating the hairstyle pairing dataset involves two main steps. First, we use the bald converter to generate bald images, which serve as proxy images for bald heads during training. Second, based on the hairstyle masks from the original images, we use a Stable Diffusion inpainting model and ChatGPT to edit the non-hairstyle portions of the original dataset, altering identities and backgrounds to create the reference images for training. Ultimately, we obtain a triplet dataset consisting of the original images, the reference images, and the bald proxy images.
Training.
In the first stage, we train the bald converter using a straightforward approach similar to ControlNet on an existing dataset. This allows us to achieve a highly effective bald converter. In the second stage, we focus on training the main components of our model. In these two training processes, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful hair transfer. These augmentations include synchronized affine transformations applied to the source image, the bald proxy image, and the target image.
The loss functions for both stages are similar to the original Stable Diffusion training objective and can be mathematically represented as follows:
$L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{s}},\mathbf{c_{r}}\right)\right\|_{2}^{2}\right],$ (1)
[…]
Inference.
During the inference process, there are two stages. In the first stage, the user-provided source images are input into the Bald Converter to obtain bald proxy images. In the second stage, the bald proxy images and the reference images are fed into our hair transfer framework to guide the generation process of diffusion.
Experiments
Figure 6: Qualitative comparison of different methods. Compared to other approaches, our method achieves more refined and stable hairstyle transfer without the need for precise facial alignment or explicit masks for supervision.
Experimental Setup
Training data.
  • FFHQ dataset subset(Karras, Laine, and Aila 2019): The FFHQ dataset is a high-quality dataset of facial images, widely used in image generation and facial recognition research. We randomly selected 20,000 images from the FFHQ dataset as the source images for our training. Using our proposed data construction method, we processed these source images to generate triplet data consisting of source images, reference images, and bald proxy images, resulting in a total of 60,000 images.
  • NH-FFHQ dataset: Non-Hair-FFHQ dataset was generated by processing the FFHQ dataset through the hairmapper(Wu, Yang, and Jin 2022) method, which is a high-quality image dataset that contains 6,000 non-hair-FFHQ portraits.
Implementation details.
We employed Stable Diffusion V1-5 as the pre-trained diffusion model. The model training process utilized a two-stage approach. In the first stage, we trained the Bald Converter using the Non-Hair FFHQ dataset. This training was conducted on a single H800 GPU with a batch size of 16 and a learning rate of 5e-5, over a total of 8,000 steps. In the second stage, we trained our Hair Extractor using the prepared triplet data. This stage was performed on 8 H800 GPUs with a batch size of 8 and a learning rate of 5e-5, over a total of 50,000 steps. During inference, we followed the same two-stage approach as used in training. Both stages utilized the DDIM sampler with 30 sampling steps and the classifier-free guidance scale setting of 1.5. The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 8 per GPU and a learning rate of 5e-5.
Evaluation Metrics
Given a hairstyle reference image and a hair color reference image, the purpose of hair transfer is to apply the corresponding hairstyle and hair color attributes to the input image. We compare our method with the current state-of-the-art methods: Barbershop(Zhu et al. 2021), SYH(Kim et al. 2022), HairFastGAN(Nikolaev et al. 2024), hairclip(Wei et al. 2022), and hairclipv2(Wei et al. 2023). All comparison algorithms use the default parameters from their official implementations.
To provide a thorough and objective assessment of the performance of each algorithm in different aspects of hairstyle transfer, we calculated the FID(Heusel et al. 2017) metrics for the source image and the generated target image. After hairstyle transfer, the structural and content information of the source image and the generated image should be consistent, so we use the SSIM(Wang et al. 2004) metric to evaluate the structural and content similarity between the source image and the generated target image. In addition, we use the PSNR metric to evaluate the quality of the generated images and we also use Insightface(Deng et al. 2019) to evaluate identity similarity(IDS) between the original source image and the generated target image.
Experiment Results
Qualitative Comparison.
As illustrated in the figure 6, we conducted qualitative comparison experiments on a variety of hairstyles. Overall, our method significantly outperforms other approaches in terms of the refinement and completeness of hairstyle transfer, while also maintaining the structural and identity consistency of the source image to a great extent. Among the methods we compared, Barbershop frequently exhibits simple hair pasting issues and often result in unnatural integration of hair and face (as seen in the third and fourth rows). Style Your Hair and Hair Fast GAN performs rather coarsely in hairstyle transfer, often neglecting hair details and showing color inconsistencies, especially with colored hair. Hairclip and Hairclip v2 demonstrate the weakest ability in hairstyle transfer, struggling to accurately transfer the reference hairstyle. In contrast, our method consistently exhibits robust and stable transfer capabilities across different hairstyle styles and colors.
Quantitative Comparison
Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
[…]
Barbershop 46.178 29.324 0.579 0.760
HairFastGAN 36.205 29.348 0.620 0.762
SYH 41.543 29.517 0.621 0.712
Hairclip 44.359 28.056 0.620 0.697
HairclipV2 37.456 30.619 0.637 0.769
Ours 33.653 29.555 0.640 0.771
Figure 7: Ablation study results of different training and inference settings.
The experiment uses the CelebA-HQ datasets as experimental data, 2500 face images are randomly selected as input from the CelebA-HQ dataset with an equal number of reference images from the remaining Celeb-HQ dataset. Table 1 shows our quantitative evaluation across different methods. Overall, our method surpasses previous approaches on most metrics. Specifically, the higher SSIM and IDS scores indicate that Stable-Hair effectively preserves the content and identity of the source image while transferring the hairstyle, demonstrating our method’s superiority in maintaining structural and content consistency. Additionally, the superior FID score suggests that our method produces high-fidelity and highly realistic results after the transfer. Our method ranks second in terms of the PSNR metric. Although it is slightly inferior to Hairclip v2 in this specific measure, our method still demonstrates high image fidelity and quality when considering other comprehensive metrics.
User Study.
Considering the subjective nature of the hairstyle transfer task, we conducted a comprehensive user study involving 30 volunteers. Specifically, we randomly sampled 20 sets of data from our quantitative experiments and selected 10 popular hairstyles from social media as reference styles, using a corresponding number of source images randomly sampled from FFHQ datasets to create an additional 10 sets of data using various algorithms. This resulted in a total of 30 triplets, each consisting of an original image, a reference image, and the transfer results. As with previous methods(Wei et al. 2022), the test results from different algorithms were randomized. For each test sample, volunteers were asked to choose the best option based on three criteria: transfer accuracy, preservation of unrelated attributes, and visual naturalness. The results shown in table 2 indicate that our method outperforms the comparison methods in transfer accuracy, preservation of unrelated attributes, and visual naturalness.
Table 2: User study on hair transfer. Accuracy denotes the accuracy for hair transfer, Preservation indicates the ability to preserve irrelevant regions and Naturalness denotes the visual realism of the generated image.
Metrics Barber Hair SYH Hair Hair Ours
shop FastGAN clip clipV2
Accuracy(%) 18.6 20.1 19.2 5.1 10.2 26.8
Preservation(%) 11.1 17.3 20.0 7.4 21.8 22.4
Naturalness(%) 13.4 20.2 15.9 11.2 18.7 20.6
Figure 8: The visual results of cross-domain transfer demonstrate the robustness of our method.
Ablation study.
To thoroughly investigate the role of each module in our method, we conducted systematic ablation studies by sequentially removing individual components. As illustrated in figure 7, it is evident that the models trained with pixel-conditioned input using ControlNet often exhibit color discrepancies (as seen in the first and third columns of the results), leading to inconsistencies between the source and target images. By employing our proposed Latent ControlNet, which maps the hair removal and transfer process from pixel space to latent space, we effectively eliminate these color inconsistencies, significantly enhancing content preservation. Additionally, the bald images generated in our Stage 1 output appear very natural, allowing for a seamless transition into Stage 2. This results in highly satisfactory transfer outcomes.
Cross-domain transfer results.
As illustrated in figure 8, the robustness of our method enables the transfer of hairstyles across diverse domains, a capability that previous approaches were unable to achieve. This demonstrates the significant advancements our method offers in the realm of hairstyle transfer.
Conclusions
In this paper, we introduce Stable-Hair, the first framework to tackle hairstyle transfer using diffusion techniques. This method marks a significant advancement, achieving stable and fine-grained real-world hairstyle transfers previously unattainable. Stable-Hair features a two-stage pipeline. The first stage uses a Bald Converter to transform the source image into a bald proxy image. The second stage comprises a Hair Extractor, a Latent IdentityNet, and Hair Cross-Attention Layers to accurately transfer the target hairstyle to the bald image. The Hair Extractor encodes reference images with the desired hairstyles, while the Latent IdentityNet preserves identity content and background consistency. The Hair Cross-Attention Layers in the U-Net ensure precise and high-fidelity hairstyle transfer. Extensive experiments demonstrate that Stable-Hair achieves commercial-grade hairstyle transfer capabilities, setting a new standard in the field.
Supplementary Materials
Preliminaries
Diffusion models.
Diffusion Model (DM) (Ho, Jain, and Abbeel 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure. Latent Diffusion Model (LDM) (Rombach et al. 2022) is proposed to model image representations in the autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
$L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$ (2)
[…]
Stable Diffusion (SD) is a widely adopted text-to-image diffusion model based on LDM. Compared to LDM, SD is trained on a large LAION (Schuhmann et al. 2022) dataset and replaces BERT with the pre-trained CLIP (Radford et al. 2021) text encoder.
Figure 9: Limitation of our method.
More Qualitative Comparison
To demonstrate the effectiveness of our method, we present more qualitative comparison results in Fig. 10. It is observed our Stable-Hair achieves the most precise and natural transfer. In contrast, all other methods failed to retain details and resulted in unnatural transfers.
Figure 10: More qualitative comparison results of hair transfer.
More Visual Results
In this section, we present additional visual results for hair transfer to demonstrate the robustness and superiority of our approach. As shown in Fig. 11 and Fig. 12, we showcase the effectiveness of our method in hair transfer on realistic domain images and hair transfer on cross-domain images.
Figure 11: More results of hair transfer for the realistic domain.
Figure 12: More results of hair transfer for cross-domain.
Broader Impact
The broader impact of diffusion-based hair style transfer methods is substantial, as they hold the potential to transform the beauty industry by facilitating more efficient and personalized hair styling applications. However, it is crucial to address the ethical considerations associated with this technology, including concerns related to privacy, consent, and the reinforcement of societal beauty standards. We explicitly discourage the unauthorized use of our method to alter the hairstyles in others’ photographs. As with any advancing technology, we advocate for a cautious approach in the deployment of diffusion-based hairstyle transfer methods and emphasize the importance of ongoing scrutiny regarding their ethical and legal ramifications.
Limitations
Due to the limitation of the training data, our method may inadvertently transfer certain hair accessories along with the makeup to the source image (As shown in Fig 9), which may not be desirable in some scenarios. We plan to address this limitation in future work to achieve more controlled and precise hair transfer.

### v2

@@
 Abstract
-Current hair transfer methods struggle to handle diverse and intricate hairstyles, thus limiting their applicability in real-world scenarios.
-In this paper, we propose a novel diffusion-based hair transfer framework, named Stable-Hair, which robustly transfers a wide range of real-world hairstyles onto user-provided faces for virtual hair try-on.
+Current hair transfer methods struggle to handle diverse and intricate hairstyles, limiting their applicability in real-world scenarios.
+In this paper, we propose a novel diffusion-based hair transfer framework, named Stable-Hair, which robustly transfers a wide range of real-world hairstyles to user-provided faces for virtual hair try-on.
 To achieve this goal, our Stable-Hair framework is designed as a two-stage pipeline.
 In the first stage, we train a Bald Converter alongside stable diffusion to remove hair from the user-provided face images, resulting in bald images.
-In the second stage, we specifically designed three modules: a Hair Extractor, a Latent IdentityNet, and Hair Cross-Attention Layers to transfer the target hairstyle with highly detailed and high-fidelity to the bald image.
-Specifically, the Hair Extractor is trained to encode reference images with the desired hairstyles.
-To preserve the consistency of identity content and background between the source images and the transfer results, we employ a Latent IdentityNet to encode the source images.
-With the assistance of our Hair Cross-Attention Layers in the U-Net, we can accurately and precisely transfer the highly detailed and high-fidelity hairstyle to the bald image.
-Extensive experiments have demonstrated that our approach delivers state-of-the-art (SOTA) results among existing hair transfer methods.
-Project page: https://xiaojiu-z.github.io/Stable-Hair.github.io/
-Figure 1: Stable-Hair is a novel diffusion-based hairstyle transfer method that can robustly transfer a diverse range of real-world hairstyles.
-We demonstrate its performance on a variety of challenging hairstyles, achieving highly detailed and high-fidelity transfers with impressive results while preserving the original identity content and structure.
+In the second stage, we specifically designed a Hair Extractor and a Latent IdentityNet to transfer the target hairstyle with highly detailed and high-fidelity to the bald image.
+The Hair Extractor is trained to encode reference images with the desired hairstyles, while the Latent IdentityNet ensures consistency in identity and background.
+To minimize color deviations between source images and transfer results, we introduce a novel Latent ControlNet architecture, which functions as both the Bald Converter and Latent IdentityNet.
+After training on our curated triplet dataset, our method accurately transfers highly detailed and high-fidelity hairstyles to the source images.
+Extensive experiments demonstrate that our approach achieves state-of-the-art performance compared to existing hair transfer methods.
+Figure 1: Stable-Hair is the first diffusion-based method for hairstyle transfer, capable of handling an extensive range of real-world hairstyles with exceptional robustness.
+Unlike previous methods, which often struggle with complex or intricate styles, Stable-Hair achieves remarkably detailed and high-fidelity transfers while preserving the original identity content.
+00footnotetext: †: Corresponding author.
 Introduction
-With the increasing prevalence of digital media and virtual reality applications, personalized virtual avatars and virtual try-on systems have emerged as a significant research area.
-Hair transfer is one of the most challenging tasks within this domain.
-The objective of this task is to transfer the hair color, shape, and structure attributes from a reference image to a user-provided source image while preserving the identity and background of the source image.
-In recent years, the advancements in Generative Adversarial Networks (GANs) (Tan et al.
+Hair transfer is one of the most challenging tasks in the virtual try-on domain.
+The objective of this task is to transfer hair color, shape, and structure attributes from a reference image to a user-provided source image while preserving the identity and background of the source image.
+In recent years, advances in Generative Adversarial Networks (GANs) (Tan et al.
 2020; Zhang and Zheng 2018; Guo et al.
 2022; Zhu et al.
-2022; Chang, Kim, and Kim ; Chung et al.
+2022b; Chang, Kim, and Kim 2023; Chung et al.
 2022; Zhu et al.
@@
 2022) have driven significant progress in this field.
-However, these GAN-based methods often struggle to handle the diverse and complex hairstyles encountered in real-world scenarios, which severely limits their effectiveness in practical applications.
-Recently, diffusion models have emerged as state-of-the-art methods in the field of image generation.
+However, these GAN-based methods often struggle to handle the diverse and complex hairstyles encountered in real-world scenarios, severely limiting their effectiveness in practical applications.
+Recently, diffusion models have emerged as SOTA methods in the field of image generation.
 These models not only enable more stable training but also demonstrate impressive results in terms of diversity and fidelity.
-Consequently, we are intrigued by the question: “Can we leverage the powerful capabilities of diffusion models to achieve more stable and high-precision hair transfer?
-”
+Consequently, we are intrigued by the question: “Can we leverage the powerful capabilities of diffusion models to achieve more stable and high-precision hair transfer?”
 Building on the aforementioned considerations, in this paper, we propose a novel hair transfer framework based on diffusion models, named Stable-Hair.
 Inspired by method (Wei et al.
-2023) which converts different editing conditions (e.g., text, sketch) into different proxies in the StyleGan W+ space, we designed a two-stage paradigm to ensure the precision and naturalness of our hair transfer process.
+2023) which converts different editing conditions (e.g., text, sketch) into different proxies in the StyleGAN $\mathcal{W}$+ space, we designed a two-stage paradigm to ensure the precision and naturalness of our hair transfer process.
 In the first stage, we train a Bald Converter to transform the user-provided source image into a bald proxy image.
-In the second stage, we utilize our specially designed model to transfer the reference hairstyle onto the bald proxy image.
-Our model is designed with three core components, each addressing specific aspects of the hair transfer challenge: the Hair Extractor, Latent IdentityNet, and Hair Cross-Attention Layers.
-The Hair Extractor is specially trained to capture and encode hairstyles from reference images with unprecedented levels of detail and texture.
-This capability enables Stable-Hair to accurately represent complex hairstyles.
-The Latent IdentityNet is used to encode the source image, ensuring that the identity and background of the source image remain consistent throughout the transfer process.
-By integrating Hair Cross-Attention Layers within diffusion U-Net, we can effectively inject the features of the reference hair into the source image.
+Furthermore, a key advancement in our methodology is the implementation of an automated data generation pipeline, which generates triplets for training.
+This pipeline utilizes the Large Language Model for generating text prompts, and the Stable Diffusion Inpainting model for creating reference images.
+This synthetic training data ensures the effective training of our framework, enabling it to robustly handle challenging hairstyles with remarkable fidelity.
+In the second stage, a Hair Extractor is specially trained to capture and encode hairstyles from reference images with unprecedented levels of detail and texture.
+A Latent IdentityNet is used to encode the source image, ensuring that the identity and background of the source image remain consistent throughout the transfer process.
+By integrating these two components with diffusion U-Net, we can effectively inject the features of the reference hair into the source image.
 This approach not only allows the hair to adapt accurately to the new environment of the source image but also ensures that it is styled appropriately, appearing natural and matching the original appearance of the subject.
-Notably, to preserve non-hair regions and the overall facial identity of the subject, we have designed a novel structure called Latent Controlnet.
+To rigorously preserve the non-hair regions and maintain the facial identity of the subject, we have developed an innovative architecture named Latent ControlNet.
 This structure integrates our Bald Converter and Latent IdentityNet.
@@
   • In this paper, we propose the first diffusion-based hair style transfer framework, named Stable-Hair.
-Compared with the GAN-based hair style transfer works, Stable-Hair achieves better fidelity and robustness.
-Experiments demonstrate that our method surpasses other state-of-the-art methods.
-  • We utilize a Hair Extractor in conjunction with Hair Cross-attention Layers to inject reference hair features.
+Experiments demonstrate that our method significantly surpasses existing SOTA hair transfer methods in terms of fidelity and fine-grained detail.
+  • We utilize a Hair Extractor to encode reference images and inject detailed hair features.
 To ensure better source content consistency during the transfer process, we introduce a novel Latent ControlNet architecture.
 This architecture is utilized as a Bald Converter and a Latent IdentityNet, which maps the hairstyle transfer process from the pixel space to the latent space.
-  • We devised an automated data production pipeline, which is a pivotal component contributing to the successful training of our framework.
+  • We utilized our designed pipeline to process both video and image data. This pipeline plays a crucial role in the effective training of our framework by generating a high-quality dataset.
 Related Works
-Hair style transfer.
+Hair Style Transfer
 The development of GAN-based methods (Tan et al.
@@
 2022; Zhu et al.
-2022; Chang, Kim, and Kim ; Chung et al.
+2022b; Chang, Kim, and Kim 2023; Chung et al.
 2022; Zhu et al.
@@
 2022) has significantly advanced the field of hairstyle transfer, with most current methods relying on GAN-based approaches.
-In the realm of hair transfer, MichiGAN(Tan et al.
-2020) decomposes hair into four orthogonal attributes and designs corresponding modules to represent, process, and convert user inputs, which are then integrated to realize an end-to-end network.
-Similarly, Barbershop(Zhu et al.
-2021) proposes a novel latent space for image blending that excels in preserving detail and encoding spatial information, extracting hairstyle information from multiple reference images.
+Figure 2: Overall schematics of our method.
+Our pipeline consists of two stages.
+First, the user’s input source image is transformed into a bald proxy image by utilizing a Bald Converter.
+In the second stage, we employ the pre-trained SD model along with a Hair Extractor to transfer the reference hair onto the bald proxy image.
+The Hair Extractor is responsible for capturing the intricate details and features of the reference hair.
+These features are then injected into the SD model through newly added hair cross-attention layers.
+After training on the triplet dataset constructed using our specially designed automated data pipeline, our method achieves highly detailed and high-fidelity hair transfers, resulting in natural and visually appealing outcomes.
+In hair transfer, MichiGAN (Tan et al.
+2020) decomposes hair into four orthogonal attributes and designs the corresponding modules to represent, process, and convert user input, which is then integrated to realize an end-to-end network.
+Barbershop (Zhu et al.
+2021) proposes a novel latent space for image blending that excels at preserving detail and encoding spatial information, extracting hairstyle information from multiple reference images.
 Furthermore, LOHO (Saha et al.
 2021) employs an optimization-based approach using GAN inversion to infill missing hair structure details in the latent space during hairstyle transfer, introducing two-stage optimization and gradient orthogonalization to enable disentangled latent space optimization of hair attributes.
-Moreover, Hairmapper(Wu, Yang, and Jin 2022) focuses on training a fully connected network to determine the direction of hair removal in the latent space of StyleGAN during the training stage.
-SYH(Kim et al.
-2022) explores a pose-aligned latent code in StyleGAN2’s latent space and proposes a high-performing pose-invariant hairstyle transfer model, which excels in transferring hairstyles under significant pose differences while preserving local hairstyle textures.
-In addition, Hairclip(Wei et al.
-2022), leveraging the powerful image-text representation capability of the Contrastive Language-Image Pre-Training (CLIP) model, proposes a unified hair editing framework based on StyleGAN.
-Building on this, HairCLIPv2(Wei et al.
-2023) converts all hair editing tasks into hair transfer tasks, with editing conditions converted into different proxies accordingly.
-It also supports a wide range of interaction modes and functions.
+Hairmapper (Wu, Yang, and Jin 2022) trains a network to direct hair removal in StyleGAN’s latent space.
+Although these early methods can transfer simple hairstyles, they do not address the challenge of pose-invariant hairstyle transfer.
+Therefore, SYH (Kim et al.
+2022) proposes a pose-invariant model that handles significant pose differences while maintaining local textures.
+Hairclip (Wei et al.
+2022) utilizes CLIP for unified hair editing, with HairCLIPv2 (Wei et al.
+2023) converting hair editing tasks into transfer tasks with varied proxies.
 StyleGAN-Salon (Khwanmuang et al.
-2023) introduces a multi-view optimization framework and the use of two guide images, representing a rough composite of the final output based on a simple cut-and-paste from the input image’s viewpoint, to optimize StyleGAN2’s learned latent space on W and W+, thereby controlling hallucinations to accurately transfer hairstyles.
+2023) employs multi-view optimization and guide images to enhance hairstyle transfer accuracy.
 HairFastGAN (Nikolaev et al.
@@
 To overcome these deficiencies, we propose the first diffusion-based method, which generates high-quality and robust images, thereby setting a new benchmark in the field.
-Diffusion models.
-Currently, diffusion models have garnered widespread attention and have seen significant advancements.
+Diffusion Models
+Currently, diffusion models have attracted a great deal of attention and have seen significant advancements.
 As the most prominent generative models today, diffusion models have achieved state-of-the-art results across various image generation tasks, including text-to-image generation (Ramesh et al.
@@
 2024c) and so on.
-Notably, the recent emergence of diffusion-based virtual try-on applications further demonstrates the formidable generative capabilities of diffusion models, enabling commercial-grade virtual try-ons (Xu et al.
+In particular, the recent emergence of diffusion-based virtual try-on applications further demonstrates the formidable generative capabilities of diffusion models, enabling commercial grade virtual try-ons (Xu et al.
 2024; Wang et al.
@@
 2024d), which were previously unattainable with traditional GAN methods.
-Diffusion models exhibit a pronounced advantage in generating high-quality images, especially when trained on extensive and diverse datasets.
-In this paper, we harness the strengths of diffusion models to produce high-fidelity and robust output images.
-This approach is particularly effective in overcoming the limitations inherent in GAN-based models, especially in the context of complex hairstyle transfer tasks.
-By leveraging the superior generative capabilities of diffusion models, our method significantly enhances the quality and realism of the transferred hairstyles, thereby addressing the challenges posed by intricate and detailed hair structures.
+Diffusion models excel in generating high-quality images, especially when trained on extensive datasets.
+In this paper, We use diffusion models to achieve high-fidelity and robust hairstyle transfers, addressing the limitations of GAN-based methods and significantly enhancing the quality and realism of complex hairstyles.
 Methodology
-Stable-Hair
-Overview.
-Figure 2: Comparative structural diagram of ControlNet and Latent ControlNet (ours).
-Our design divides the hairstyle transfer process into two stages.
+Overview
+As shown in Fig.
+2, our design divides the hairstyle transfer process into two stages.
 Firstly, the user’s input source image is transformed into a bald proxy image by using the Bald Converter.
-Secondly, our model is employed to transfer the reference hair onto the bald proxy image.
-This ensures optimal stability in hairstyle transfer and maintains the source image content.
-Figure 3: Overall schematics of our method.
-Our pipeline consists of two stages to achieve high-quality hair transfer.
-First, the user’s input source image is transformed into a bald proxy image.
-This transformation is accomplished using a pre-trained Stable Diffusion (SD) model in conjunction with a specialized Bald Converter.
-In the second stage, we employ the pre-trained SD model along with a Hair Extractor to transfer the reference hair onto the bald proxy image.
-The Hair Extractor is responsible for capturing the intricate details and features of the reference hair.
-These features are then injected into the SD model through newly added hair cross-attention layers.
-By leveraging these two stages, our method achieves highly detailed and high-fidelity hair transfers, producing natural and visually appealing results.
-As shown in Fig 3, our Stable-Hair framework comprises three primary components: Hair Extractor, Latent IdentityNet, and the newly integrated Hair Cross-Attention Layers.
-Given a reference image $\mathit{I_{r}}$ containing the desired hairstyle and a source image $\mathit{I_{s}}$ provided by users, Hair Extractor can stably and precisely encode a variety of reference hairstyles in real-world scenes and extract detailed hair features $\mathit{F_{h}}$.
-These detailed hair features $\mathit{F_{h}}$ are fed to the diffusion U-Net structure through Hair Cross-Attention Layers.
-The Latent IdentityNet modules are employed for processing bald proxy image $\mathit{I_{b}}$ which is converted from the source image $\mathit{I_{s}}$ by Bald Converter.
-Figure 4: Accumulated color deviations in different designing options.
-Latent ControlNet.
+Secondly, our model is used to transfer the reference hair onto the bald proxy image.
+This two-stage approach ensures optimal stability in hairstyle transfer and maintains the source image content.
+Specifically, given a reference image containing the desired hairstyle and a source image provided by users, Hair Extractor can stably and precisely encode a variety of reference hairstyles in real-world scenes and extract detailed hair features.
+These detailed hair features are fed to the diffusion U-Net structure.
+The Latent IdentityNet modules are used to process the bald proxy image that is converted from the source image by the Bald Converter.
+Stage1: Standardizing to Bald State
+Bald Converter.
 In our method, the source image is processed in two steps, the first step is to convert the source image into a bald proxy image by a Bald Converter.
-The second step is to maintain the content of the non-hair region of the bald proxy image using Latent IdentityNet.
-The maintenance of content consistency in the source image is crucial through these two steps.
-Any deviation from the intended goal, such as alterations in color or structure, would result in a content-inconsistent final image.
-Therefore, designing these two content maintenance modules becomes a pivotal aspect of our hair transfer framework.
-A naive baseline approach involves utilizing the ControlNet structure as a Bald Converter and Latent IdentityNet to ensure content consistency.
-However, our experimental findings indicate that while ControlNet effectively maintains the structural consistency of the source image, it struggles to preserve color consistency.
+Specifically, we use the vanilla diffusion model in conjunction with our designed Latent ControlNet structure (introduced in the following section) to transfer the source image into the bald state.
+After training, the Bald Converter achieves complete hair removal without requiring image alignment or cropping.
+This results in a cleaner, more consistent baseline for subsequent hairstyle transfer, leading to improved visual fidelity and more realistic outcomes.
+Notably, standardizing source images to a bald state is crucial for enhancing model performance in hairstyle transfer.
+This process allows the model to concentrate more effectively on facial features and details by removing hair-related variability.
+Such standardization is instrumental in improving both training stability and transfer accuracy.
+The significance and impact of this approach are discussed in greater detail in the supplementary materials.
+Figure 3: Synthetic Training Data: We propose an automated data generation pipeline to generate {Original image (or frame 1), Reference image, Bald proxy image} triplets for training.
+The pipeline uses ChatGPT to generate text prompts, the Stable Diffusion Inpainting model to generate reference images, and our pre-trained Bald converter to convert the original image or one of the frames sampled from videos into the bald proxy image.
+Training Data Collection
+After training our Bald Converter, we can utilize the Bald Converter alongside the Stable Diffusion Inpainting model and ChatGPT to generate the triplet training data necessary to train the second-stage models.
+Automatic Pipeline.
 As shown in Fig.
-4, due to accumulated color deviations over these two steps, there are noticeable changes in final colors.
-Why does the controlnet have a color difference?
-We supposed that the reason is that pixel space and latent space represent image information in fundamentally different ways.
-Pixel space deals with the raw pixel values of an image, while latent space involves a more abstract, high-dimensional representation created by VAE encoder.
-For diffusion models, the characteristics and distribution of information in these two spaces can vary significantly, making alignment challenging during training.
-Therefore, as depicted in Fig.
-2, we improve the ControlNet structure and propose a new variant named Latent ControlNet.
-Before the image is input into the ControlNet, the image is encoded into the latent space by the VAE encoder and then sent to the trainable copy of U-Net after a new trainable convolutional layer.
-Finally, we train our Bald Converter and Latent IdentityNet based on our proposed Latent ControlNet structure and get the best consistency practice.
-Hair Extractor.
-Hair transfer needs to transfer the hair on the reference image $\mathit{I_{r}}$ in a detailed, complete, and accurate way, and our Hair Extraxtor is designed to achieve such transfer.
-Inspired by recent works (Wang et al.
-2024; Zhang et al.
-2024a) on reference image-guided generation, we utilize a trainable copy of U-Net architecture from pre-trained diffusion models as our Hair Extractor.
-Specifically, we encode our reference image $\mathit{I_{r}}$ through Hair Extractor and collect the features of the self-attention layer in each transformer block as detailed hair features $\mathit{F_{h}}$.
-Then the features are subsequently injected into diffusion U-Net through Hair Cross-Attention Layers.
-Hair Cross-Attention Layers.
-The goal of the Hair Cross-Attention Layers is to integrate detailed hair features into the U-Net and align them appropriately with the intermediate feature maps of U-Net, enabling the model to accurately transfer the hair onto the appropriate region in the source image.
-Specifically, as shown in the figure, in each transformer block of the U-Net, we retain the original self-attention layers and add Hair Cross-Attention Layers.
-The detailed Hair Features $\mathit{F_{h}}$ are input to the hair cross-attention layers and serve as the K(key) and V(value) features.
-The hair cross-attention layers and the original self-attention layers in the U-Net share the Q feature.
-With the help of the cross-attention mechanism, the hair and source face are well-aligned.
-Finally, we simply add the output of hair cross-attention to the output of original self-attention.
-Figure 5: Synthetic Training Data: We propose an automated data generation pipeline to generate (Original image, Reference image, Bald proxy image) triplets for training.
-The pipeline uses ChatGPT to generate text prompts, the Stable Diffusion Inpainting model to generate reference images, and our pre-trained Bald converter to convert the original image into the bald proxy image.
-Model Training and Inference
-Training Data Collection.
-As shown in Fig.
-5 Our pipeline for creating the hairstyle pairing dataset involves two main steps.
+3, our pipeline to create the hairstyle pairing dataset involves two main steps.
 First, we use the bald converter to generate bald images, which serve as proxy images for bald heads during training.
@@
 Ultimately, we obtain a triplet dataset consisting of the original images, the reference images, and the bald proxy images.
-Training.
+Video Strategy.
+Relying solely on image data, the reference and source images typically share the same head pose.
+This limitation prevents the model from effectively handling cases where the head poses in the reference and source images are misaligned.
+Video data, on the other hand, often includes frames of the same individual (with consistent hairstyle) from different angles.
+By using video data to generate our dataset—selecting one frame as the source image and another with a different pose as the reference image—we can elegantly address this issue.
+This approach enhances the model’s robustness to pose variations.
+In terms of image data, we used every single image as original images to generate reference images and bald proxy images, resulting in a total of 60,000 images.
+For video data, we sampled two frames from the same video.
+One frame was used to generate bald proxy images, and another frame was used to generate reference images.
+Resulting in a total of 90,000 images.
+Stage2: Hair Transfer and Integration
+Latent IdentityNet.
+In our method, the second step is to transfer the reference hair onto the bald proxy image.
+The maintenance of content consistency in the source image is crucial through our two steps.
+Any deviation from the intended goal, such as alterations in color or identity, would result in a content-inconsistent final image.
+Therefore, designing a maintenance module becomes a pivotal aspect of our hair transfer framework.
+A baseline approach is using the ControlNet structure as a Bald Converter and Latent IdentityNet to ensure content consistency.
+However, our experimental findings indicate that while ControlNet effectively maintains the structural consistency of the source image, it struggles to preserve color consistency.
+As shown in Fig.
+6, due to accumulated color deviations over these two steps, there are noticeable changes in the final colors.
+Why does the controlnet have a color difference?
+We assumed that the reason is that the pixel space in ControlNet and the latent space in U-Net represent image information in fundamentally different ways.
+The pixel space deals with the raw pixel values of an image, while the latent space involves a more abstract, high-dimensional representation created by the VAE encoder.
+For diffusion models, the misalignment between the characteristics and distribution of information in these two spaces can lead to difficulties in maintaining color consistency during the hair transfer process.
+Therefore, as shown in Fig.
+2 (right), we improve the ControlNet structure and propose a new variant named Latent ControlNet.
+Before the image is input into the ControlNet, the image is encoded into the latent space by the VAE encoder and then sent to the trainable copy of U-Net after a new trainable convolutional layer.
+Finally, we train our Bald Converter and Latent IdentityNet based on our proposed Latent ControlNet structure and get the best consistency practice.
+Figure 4: Qualitative comparison of different methods.
+Compared to other approaches, our method achieves more refined and stable hairstyle transfer without the need for precise facial alignment or explicit masks for supervision.
+Hair Extractor.
+Hair transfer needs to transfer the hair in the reference image in a detailed, complete, and accurate way, and our Hair Extraxtor is designed to achieve such transfer.
+Inspired by recent works (Wang et al.
+2024; Zhang et al.
+2024a) on reference image-guided generation, we utilize a trainable copy of the U-Net from pre-trained diffusion models as our Hair Extractor.
+Specifically, we encode our reference image through the Hair Extractor and collect the features of the self-attention layer in each transformer block as detailed hair features.
+The features are then subsequently injected into the diffusion U-Net through newly added hair cross-attention layers.
+In each transformer block of the U-Net, we retain the original self-attention layers and add hair cross-attention layers.
+Detailed hair features are entered into the hair cross-attention layers and serve as the K(key) and V(value) features.
+Both attention layers in the U-Net share the Q feature.
+Finally, we simply add the output of hair cross-attention to the output of original self-attention.
+Model Training
 In the first stage, we train the bald converter using a straightforward approach similar to ControlNet on an existing dataset.
 This allows us to achieve a highly effective bald converter.
-In the second stage, we focus on training the main components of our model.
+In the second stage, we focus on training the Hair Extractor and the Latent IdentityNet.
 In these two training processes, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful hair transfer.
 These augmentations include synchronized affine transformations applied to the source image, the bald proxy image, and the target image.
-The loss functions for both stages are similar to the original Stable Diffusion training objective and can be mathematically represented as follows:
+The loss functions for both stages can be mathematically represented as follows:
 $L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{s}},\mathbf{c_{r}}\right)\right\|_{2}^{2}\right],$ (1)
 […]
-Inference.
-During the inference process, there are two stages.
-In the first stage, the user-provided source images are input into the Bald Converter to obtain bald proxy images.
-In the second stage, the bald proxy images and the reference images are fed into our hair transfer framework to guide the generation process of diffusion.
 Experiments
-Figure 6: Qualitative comparison of different methods.
-Compared to other approaches, our method achieves more refined and stable hairstyle transfer without the need for precise facial alignment or explicit masks for supervision.
-Experimental Setup
-Training data.
-  • FFHQ dataset subset(Karras, Laine, and Aila 2019): The FFHQ dataset is a high-quality dataset of facial images, widely used in image generation and facial recognition research.
-We randomly selected 20,000 images from the FFHQ dataset as the source images for our training.
-Using our proposed data construction method, we processed these source images to generate triplet data consisting of source images, reference images, and bald proxy images, resulting in a total of 60,000 images.
-  • NH-FFHQ dataset: Non-Hair-FFHQ dataset was generated by processing the FFHQ dataset through the hairmapper(Wu, Yang, and Jin 2022) method, which is a high-quality image dataset that contains 6,000 non-hair-FFHQ portraits.
-Implementation details.
+Implementation Details
 We employed Stable Diffusion V1-5 as the pre-trained diffusion model.
 The model training process utilized a two-stage approach.
-In the first stage, we trained the Bald Converter using the Non-Hair FFHQ dataset.
+In the first stage, we trained the Bald Converter using the Non-Hair FFHQ dataset (Wu, Yang, and Jin 2022).
 This training was conducted on a single H800 GPU with a batch size of 16 and a learning rate of 5e-5, over a total of 8,000 steps.
-In the second stage, we trained our Hair Extractor using the prepared triplet data.
-This stage was performed on 8 H800 GPUs with a batch size of 8 and a learning rate of 5e-5, over a total of 50,000 steps.
+In the second stage, we trained our Hair Extractor and Latent IdentityNet using the prepared triplet data.
+This stage was performed on 8 H800 GPUs with a batch size of 8 and a learning rate of 5e-5, over a total of 100,000 steps.
 During inference, we followed the same two-stage approach as used in training.
 Both stages utilized the DDIM sampler with 30 sampling steps and the classifier-free guidance scale setting of 1.5.
-The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 8 per GPU and a learning rate of 5e-5.
+Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
+Method CLIP-I $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ IDS $\uparrow$
+Barbershop 0.431 46.178 30.397 0.629 0.760
+HairFastGAN 0.426 36.205 30.383 0.666 0.762
+SYH 0.429 41.543 30.489 0.645 0.712
+Hairclip 0.391 44.359 28.855 0.615 0.697
+Hairclip v2 0.419 37.456 31.552 0.642 0.769
+[…]
 Evaluation Metrics
-Given a hairstyle reference image and a hair color reference image, the purpose of hair transfer is to apply the corresponding hairstyle and hair color attributes to the input image.
-We compare our method with the current state-of-the-art methods: Barbershop(Zhu et al.
+Given a hairstyle reference image, the purpose of hair transfer is to apply the corresponding hairstyle and hair color attributes to the input image.
+We compare our method with current state-of-the-art methods: Barbershop (Zhu et al.
 2021), SYH (Kim et al.
@@
 2017) metrics for the source image and the generated target image.
-After hairstyle transfer, the structural and content information of the source image and the generated image should be consistent, so we use the SSIM(Wang et al.
-2004) metric to evaluate the structural and content similarity between the source image and the generated target image.
-In addition, we use the PSNR metric to evaluate the quality of the generated images and we also use Insightface(Deng et al.
+After hairstyle transfer, the identity and background information of the source image and the generated image should be consistent, so we use SSIM (Wang et al.
+2004) and PSNR to evaluate the identity and background similarity between the source image and the generated target image.
+Notably, PSNR and SSIM are calculated at the intersected non-hair regions before and after editing.
+We also use Insightface (Deng et al.
 2019) to evaluate identity similarity (IDS) between the original source image and the generated target image.
+To evaluate hairstyle transfer capabilities, we use CLIP-I(Radford et al.
+2021) as a metric, which calculates the cosine similarity between the image embeddings of the reference image and the transferred hairstyle image.
 Experiment Results
 Qualitative Comparison.
-As illustrated in the figure 6, we conducted qualitative comparison experiments on a variety of hairstyles.
-Overall, our method significantly outperforms other approaches in terms of the refinement and completeness of hairstyle transfer, while also maintaining the structural and identity consistency of the source image to a great extent.
-Among the methods we compared, Barbershop frequently exhibits simple hair pasting issues and often result in unnatural integration of hair and face (as seen in the third and fourth rows).
-Style Your Hair and Hair Fast GAN performs rather coarsely in hairstyle transfer, often neglecting hair details and showing color inconsistencies, especially with colored hair.
+As illustrated in Fig.
+4, we conducted qualitative comparison experiments on a variety of hairstyles.
+Overall, our method significantly outperforms other approaches in terms of the refinement and completeness of hairstyle transfer, while also maintaining the background and identity consistency of the source image to a great extent.
+Among the methods we compared, Barbershop tends to produce chaotic results for complex hairstyle transfers, as observed in the second and fifth rows.
+Style Your Hair and HairFastGAN perform relatively coarsely in hairstyle transfer, often neglecting the fine details of hair texture and color.
 Hairclip and Hairclip v2 demonstrate the weakest ability in hairstyle transfer, struggling to accurately transfer the reference hairstyle.
 In contrast, our method consistently exhibits robust and stable transfer capabilities across different hairstyle styles and colors.
-Quantitative Comparison
-Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
-[…]
-Barbershop 46.178 29.324 0.579 0.760
-HairFastGAN 36.205 29.348 0.620 0.762
-SYH 41.543 29.517 0.621 0.712
-Hairclip 44.359 28.056 0.620 0.697
-HairclipV2 37.456 30.619 0.637 0.769
-Ours 33.653 29.555 0.640 0.771
-Figure 7: Ablation study results of different training and inference settings.
-The experiment uses the CelebA-HQ datasets as experimental data, 2500 face images are randomly selected as input from the CelebA-HQ dataset with an equal number of reference images from the remaining Celeb-HQ dataset.
+Furthermore, We also compared the hair removal capabilities of our Bald Converter with HairMapper in Fig.
+5.
+Our Bald Converter demonstrates robust and reliable performance, effectively transforming source images across various challenging scenarios.
+In contrast, HairMapper frequently struggles in these contexts, leading to inconsistencies in both identity and background.
+Quantitative Comparison.
+Figure 5: Visual comparison of hair removal using HairMapper.
+Our bald converter demonstrates robust performance, effectively converting source images across a wide range of poses, half-body shots, and even animated characters.
+In contrast, HairMapper struggles with these diverse scenarios, failing to maintain identity and background consistency.
+The experiment uses the CelebA-HQ dataset (Karras et al.
+2017) as experimental data, 2500 face images are randomly selected as input from the CelebA-HQ dataset with an equal number of reference images from the remaining Celeb-HQ dataset.
 Table 1 shows our quantitative evaluation across different methods.
-Overall, our method surpasses previous approaches on most metrics.
-Specifically, the higher SSIM and IDS scores indicate that Stable-Hair effectively preserves the content and identity of the source image while transferring the hairstyle, demonstrating our method’s superiority in maintaining structural and content consistency.
-Additionally, the superior FID score suggests that our method produces high-fidelity and highly realistic results after the transfer.
-Our method ranks second in terms of the PSNR metric.
-Although it is slightly inferior to Hairclip v2 in this specific measure, our method still demonstrates high image fidelity and quality when considering other comprehensive metrics.
+Overall, our method outperforms previous approaches on the majority of metrics, attaining either the top or second position across all evaluated criteria.
+Specifically, the highest SSIM and IDS scores, along with a notable PSNR ranking, demonstrate that Stable-Hair effectively preserves both the background and identity of the source image during hairstyle transfer.
+Additionally, the lowest FID score indicates that our method achieves superior visual fidelity and realism in the transferred hairstyles.
+Achieving the highest CLIP-I score indicates that our method excels at aligning the transferred hairstyle with the reference image, showing strong performance in maintaining the desired style.
 User Study.
@@
 The results shown in Table 3 indicate that our method outperforms the comparison methods in transfer accuracy, preservation of unrelated attributes, and visual naturalness.
+Table 2: Quantitative ablation of different design options. (‘pixel’ refers to the use of the original ControlNet, while ’latent’ denotes the use of the Latent ControlNet.)
+[…]
+Ours (latent) 0.434 35.128 30.980 0.680 0.779
+Ours (pixel) 0.423 39.236 29.317 0.668 0.770
+Figure 6: Visual ablation study results comparing different design options.
+The Latent ControlNet clearly maintains color consistency, whereas the original ControlNet introduces color discrepancies at each stage.
+Ablation Study.
+To thoroughly investigate the role of each module in our method, we conducted systematic ablation.
+As illustrated in Fig.
+6, it is evident that models trained with pixel-conditioned input using ControlNet often exhibit color discrepancies (as seen in the first and third columns of the results), leading to inconsistencies between the source and target images.
+Using our proposed Latent ControlNet, which maps the hair removal and transfer process from pixel space to latent space, we effectively eliminate these color inconsistencies, significantly enhancing content preservation.
+The results of the quantitative ablation are shown in Table 4.
+Our method consistently outperforms across all metrics, highlighting the superior transfer capabilities and the enhanced preservation of identity and background achieved by the Latent ControlNet.
+To further investigate the impact of video data on our method, we conducted training using only image data while keeping the same parameter settings.
+The results of the visual ablation study, as shown in Fig.
+7, clearly indicate that when trained solely with image data, our method struggles to handle significant variations in facial poses from the source images, leading to a collage-like effect with less natural transitions.
+In contrast, when trained with a mix of video data, our method demonstrates robustness to changes in facial pose.
+See the supplements for additional ablation studies.
 Table 3: User study on hair transfer.
 Accuracy denotes the accuracy for hair transfer, Preservation indicates the ability to preserve irrelevant regions and Naturalness denotes the visual realism of the generated image.
+Our method achieves the best results in all three categories.
 Metrics Barber Hair SYH Hair Hair Ours
@@
 Naturalness(%) 13.4 20.2 15.9 11.2 18.7 20.6
-Figure 8: The visual results of cross-domain transfer demonstrate the robustness of our method.
-Ablation study.
-To thoroughly investigate the role of each module in our method, we conducted systematic ablation studies by sequentially removing individual components.
-As illustrated in figure 7, it is evident that the models trained with pixel-conditioned input using ControlNet often exhibit color discrepancies (as seen in the first and third columns of the results), leading to inconsistencies between the source and target images.
-By employing our proposed Latent ControlNet, which maps the hair removal and transfer process from pixel space to latent space, we effectively eliminate these color inconsistencies, significantly enhancing content preservation.
-Additionally, the bald images generated in our Stage 1 output appear very natural, allowing for a seamless transition into Stage 2.
-This results in highly satisfactory transfer outcomes.
-Cross-domain transfer results.
-As illustrated in figure 8, the robustness of our method enables the transfer of hairstyles across diverse domains, a capability that previous approaches were unable to achieve.
-This demonstrates the significant advancements our method offers in the realm of hairstyle transfer.
+Figure 7: Visual ablation results of different training data.
+When trained exclusively on image data, the model often resorts to a simplistic copy-paste of the hairstyle.
+However, by integrating video data into the training process, our method gains pose robustness, enabling adaptive and natural hair transfer based on facial orientation.
 Conclusions
@@
 The first stage uses a Bald Converter to transform the source image into a bald proxy image.
-The second stage comprises a Hair Extractor, a Latent IdentityNet, and Hair Cross-Attention Layers to accurately transfer the target hairstyle to the bald image.
-The Hair Extractor encodes reference images with the desired hairstyles, while the Latent IdentityNet preserves identity content and background consistency.
-The Hair Cross-Attention Layers in the U-Net ensure precise and high-fidelity hairstyle transfer.
+Following this step, we designed an automated data collection pipeline that leverages the Bald Converter, an inpainting model, and ChatGPT to gather a diverse and large-scale dataset of triplets for training the second stage.
+In the second stage, a Hair Extractor and a Latent IdentityNet are utilized to accurately and robustly transfer the target hairstyle onto the bald image.
 Extensive experiments demonstrate that Stable-Hair achieves commercial-grade hairstyle transfer capabilities, setting a new standard in the field.
@@
 Diffusion Models.
-Diffusion Model (DM) (Ho, Jain, and Abbeel 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure.
+Diffusion Model (DM) (Ho, Jain, and Abbeel 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to the target data distribution $\mathbf{x_{0}}$ using an iterative denoising procedure.
 Latent Diffusion Model (LDM) (Rombach et al.
@@
 LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions.
-The LDM loss is:
+The LDM loss is as follows:
 $L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$ (2)
@@
 2021) text encoder.
-Figure 9: Limitation of our method.
+Source Data
+We utilize these three datasets to collect our training data.
+  • FFHQ dataset subset (Karras, Laine, and Aila 2019): We randomly selected 20,000 images from the FFHQ dataset as the original image data for our training.
+  • CelebV-HQ (Zhu et al. 2022a): Large-scale, high-quality video data containing 35,666 clips involving 15,653 identities and 83 manually labeled facial attributes.
+  • NH-FFHQ dataset: Non-Hair-FFHQ dataset was generated by processing the FFHQ dataset through the hairmapper (Wu, Yang, and Jin 2022) method, which is a high-quality image dataset that contains 6,000 non-hair-FFHQ portraits.
+Table 4: Quantitative ablation of different training data.
+[…]
+Ours (image+video) 0.434 35.128 30.980 0.680 0.779
+Ours (only image) 0.431 33.653 30.541 0.676 0.771
+Quantitative Ablation of Training Data
+In this section, we further examine the quantitative impact of incorporating video data into the training set on the performance of hairstyle transfer.
+As shown in Table 4, when the training data consists solely of images, the FID metric shows improvement, while other metrics exhibit slight declines compared to the case where the full dataset is used.
+This indicates that while there is a slight decline in fidelity and overall image generation quality when video data is incorporated into the training, the model’s ability to transfer hairstyles and maintain identity and background consistency is enhanced
+Why Use Bald Converter?
+First, removing the original hairstyle eliminates the need for the model to blend the new hairstyle with any remnants of the old one, reducing visual artifacts and resulting in a more natural and visually appealing transfer.
+Secondly, by standardizing the source images to a bald state, the model can better focus on facial features and structure, enhancing the accuracy and realism of the hairstyle transfer.
+Additionally, this preprocessing step helps to reduce variability in the training data, improving the model’s generalization capability across different hairstyles.
+Lastly, converting to a bald state simplifies the alignment process, ensuring that the new hairstyle seamlessly integrates with the subject’s face, even in cases of diverse facial poses.
+To further demonstrate the importance of standardizing the source image’s hairstyle to a bald state in Stage 1, we conducted a set of comparative experiments.
+In these experiments, rather than standardizing to a bald state as in Stage 1, we randomly edited the source image’s hairstyle using an inpainting model and then trained a hair transfer model based on these edited images.
+The training framework and data references are illustrated in the figure 9, with all training settings consistent with those described in our experiments section.
+The visual comparison results are also shown in the figure 10.
+As evident from the results, our model, which standardizes the hairstyle to a bald state, achieves high-fidelity and precise hair transfer, whereas the model trained on randomly edited hairstyles fails to exhibit any effective hairstyle transfer capability.
+Additionally, our method achieves hair removal without requiring image alignment or cropping, ensuring that the entire hairstyle is accurately removed and no residual hair remains.
+This guarantees a cleaner, more consistent starting point for subsequent hairstyle transfer, leading to improved visual fidelity and more realistic results.
+As shown in Fig.
+12, our method is capable of removing hair from images with various poses, half-body shots, and even cartoon or animated characters.
 More Qualitative Comparison
 To demonstrate the effectiveness of our method, we present more qualitative comparison results in Fig.
-10.
+11.
 It is observed our Stable-Hair achieves the most precise and natural transfer.
 In contrast, all other methods failed to retain details and resulted in unnatural transfers.
-Figure 10: More qualitative comparison results of hair transfer.
-More Visual Results
+More Transfer Results
 In this section, we present additional visual results for hair transfer to demonstrate the robustness and superiority of our approach.
@@
 13 and Fig.
-12, we showcase the effectiveness of our method in hair transfer on realistic domain images and hair transfer on cross-domain images.
-Figure 11: More results of hair transfer for the realistic domain.
-Figure 12: More results of hair transfer for cross-domain.
+14, we showcase the effectiveness of our method in hair transfer on realistic domain images and hair transfer on cross-domain images.
+Figure 8: Limitation of our method.
+Limitations
+Due to the limitation of the training data, our method may inadvertently transfer certain hair accessories to the source image (as shown in Fig.
+8), which may not be desirable in some scenarios.
+We plan to address this limitation in future work to achieve more controlled and precise hair transfer.
 Broader Impact
@@
 As with any advancing technology, we advocate for a cautious approach in the deployment of diffusion-based hairstyle transfer methods and emphasize the importance of ongoing scrutiny regarding their ethical and legal ramifications.
-Limitations
-Due to the limitation of the training data, our method may inadvertently transfer certain hair accessories along with the makeup to the source image (As shown in Fig 9), which may not be desirable in some scenarios.
-We plan to address this limitation in future work to achieve more controlled and precise hair transfer.
+Figure 9: The training framework without the bald converter. (i.e., the source data is not standardized to the bald state)
+Figure 10: Visual comparison of the results under different data strategies reveals significant differences.
+When the bald converter is not applied (i.e., the source data is not standardized to the bald state), the model fails to learn how to transfer hairstyles during training and instead generates random hairstyles.
+However, when the data is standardized to the bald state, the model incrementally learns to transfer high-fidelity and high-precision hairstyles.
+Figure 11: More qualitative comparison results of hair transfer.
+Figure 12: More visual results of hair removal.
+Figure 13: More results of hair transfer for the realistic domain.
+Figure 14: More results of hair transfer for cross-domain.